Let{N(t),t≥0}beaPoissonprocesswithrateλ.Fori≤nands<t,

(a) findP(N(t)=n|N(s)=i);

(b) findP(N(s)=i|N(t)=n).
Answer #1

Solidon : compede the condidonal a het oesat clisdobudós N(5) lact) =D P(NES) = INCE)=0) =P ( N(S)-KNCH)=) PONCE=) DIN(0,5) =P(N(9) = k ln(t)=n] N B(0, 514). K=0,1, 20 924 b) NCED=0 | N(S)= ;) = (NCE) = 0, n(s)=;). P() = ;) - P (NCS) - ;) PINCH-5) =